OpenStudy (anonymous):

Petrolyn motor oil is a combination of natural oil and synthetic oil. It contains 4 liters of natural oil for every 7 liters of synthetic oil. If 399 liters of synthetic oil are used, how many liters of Petrolyn oil will be made?

OpenStudy (anonymous):

@Hero, can you please help me again?

OpenStudy (anonymous):

I mean 57*4= x x+399=answer

hero (hero):

The ratio of natural oil to synethic oil is \[\frac{4}{7}\] \[\frac{4}{7} = \frac{x}{399}\] Solve that to find how much natural oil there is. Then 399 + x = the amount of Petrolyn oil used.